OpenStudy (anonymous):

Arthur and Betty start walking toward each other when they are 100 m apart. Arthur has a speed of 3.0 m/s and Betty has a speed of 2.0 m/s. Their dog, Spot, starts by Arthur's side at the same time and runs back and forth between them at 5.0 m/s. By the time Arthur and Betty meet, what distance has Spot run?

OpenStudy (anonymous):

 Suppose, Arthur and Betty meet at the point C. We know that: \(\Large S_A=v_At\) \(\Large S_B=-v_Bt\) When they meet, \(\Large S_A+S_B=L \Leftrightarrow v_At - v_Bt=L\) \(\Large 3t-2t=100 \Rightarrow t=100\) \(\large s\) So, Spot run distance: \(\Large S_S=v_St=5\times 100=500\) \(\large m\)